# Learning Decision-Focused Uncertainty Sets in Robust Optimization

### v4

@@ E.4 Polyhedral reformulation
$\begin{array}[t]{ll}\underset{Az+b\in S}{\sup}&g(x,Az+b)\\ \mbox{subject to}&Dz\leq d\end{array}$
$=\begin{array}[t]{ll}\underset{Az+b\in S}{\sup}~{}\underset{h\geq 0}{\inf}~{}g(x,Az+b)-(Dz)^{T}h+d^{T}h\\ \end{array}$
@@
$=\begin{array}[t]{ll}\underset{h\geq 0,\tau}{\inf}&\tau\\ \mbox{subject to}&d^{T}h_{l}+\underset{Az+b\in S}{\sup}g_{l}(x,Az+b)-(Dz)^{T}h_{l}\leq\tau,\quad l=1,\dots,L\end{array}$
$=\begin{array}[t]{ll}\underset{h\geq 0,\omega,\tau,y,\gamma}{\inf}&\tau\\ \mbox{subject to}&d^{T}h_{l}+[-g_{l}]^{\star}(x,\omega_{l})-b^{T}(\omega_{l}+\gamma_{l})+\sigma_{S}(\gamma_{l})+\mathcal{I}_{D^{T}h_{l}=-y_{l}}\\ &\hfill+\mathcal{I}_{A^{T}(\omega_{l}+\gamma_{l})=y_{l}}\leq\tau,\quad l=1,\dots,L\end{array}$
$=\begin{array}[t]{ll}\underset{z\geq 0,\omega,\tau,\gamma}{\inf}&\tau\\ \mbox{subject to}&[-g_{l}]^{\star}(x,\omega_{l})-b^{T}(\omega_{l}+\gamma_{l})+\sigma_{S}(\gamma_{l})\leq\tau,\quad l=1,\dots,L\\ &A^{T}(\omega_{l}+\gamma_{l})=-D^{T}h_{l},\quad l=1,\dots,L,\end{array}$
@@
$\begin{array}[t]{ll}{\text{minimize}}&f(x)\\ \mbox{subject to}&[-g_{l}]^{\star}(x,\omega_{l})-b^{T}(\omega_{l}+\gamma_{l})+\sigma_{S}(\gamma_{l})\leq 0,\quad l=1,\dots,L\\ &A^{T}(\omega_{l}+\gamma_{l})=-D^{T}h_{l},\quad l=1,\dots,L.\end{array}$
@@ E.5 Mean Robust Optimization reformulation
We adapt the proof from (Wang et al.,, 2022, Appendix A.3). To simplify notation, we define $c_{k}(v_{lk})\coloneqq\|v_{lk}-\bar{d}_{k}\|^{p}-\epsilon^{p}$. We also note that $u_{lk}=Av_{lk}+b$. Then, we have
$\begin{array}[t]{ll}\underset{u_{11},\dots,u_{LK}\in S,\alpha\in\Gamma}{\sup}&\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}g_{l}(x,u_{lk})\\ \mbox{~{}~{}~{}~{}subject to}&\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}c_{k}(v_{lk})\leq 0\end{array}$
$=\begin{array}[t]{ll}\underset{u_{11},\dots,u_{LK}\in S,\alpha\in\Gamma}{\sup}&\underset{\lambda\geq 0}{\inf}~{}\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}g_{l}(x,u_{lk})-\lambda\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}c_{k}(v_{lk})\\ \end{array}$
$=\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0}{\inf}~{}\underset{u_{11},\dots,u_{LK}\in S}{\sup}&\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}g_{l}(x,u_{lk})-\lambda\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}c_{k}(v_{lk}).\\ \end{array}$
@@
$=\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0}{\inf}&\sum_{k=1}^{K}s_{k}\\ \mbox{subject to}~{}\underset{u_{11},\dots,u_{LK}\in S}{\sup}&\sum_{l=1}^{L}\alpha_{lk}(g_{l}(x,u_{lk})-\lambda c_{k}(v_{lk}))\leq s_{k},\quad k=1,\dots,K\end{array}$
$=\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0}{\inf}&\sum_{k=1}^{K}s_{k}\\ \mbox{subject to}~{}\underset{\alpha_{11}u_{11},\dots,\alpha_{LK}u_{LK}\in S^{\prime}}{\sup}&\sum_{l=1}^{L}\alpha_{lk}(g_{l}(x,(\alpha_{lk}u_{lk})/\alpha_{lk})\\ &-\lambda c_{k}((\alpha_{lk}v_{lk})/\alpha_{lk}))\leq s_{k},\quad k=1,\dots,K.\end{array}$
In the last step, we rewrote $u_{lk}=(\alpha_{lk}u_{lk})/\alpha_{lk}$, $v_{lk}=(\alpha_{lk}v_{lk})/\alpha_{lk}$, and maximized over $\alpha_{lk}u_{lk}\in S^{\prime}$, where $S^{\prime}$ is the support transformed by $\alpha_{lk}$. See Wang et al., (2022) for details. Next, we make substitutions $h_{ij}=\alpha_{lk}u_{lk}$, $z_{ij}=\alpha_{lk}v_{lk}$, and define functions $g_{l}^{\prime}(x,h_{lk})=\alpha_{lk}g_{l}(x,h_{lk}/\alpha_{lk})$, $c_{k}^{\prime}(z_{lk})=\alpha_{lk}c_{k}^{\prime}(z_{lk}/\alpha_{lk})$.
$=\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0}{\inf}&\sum_{k=1}^{K}s_{k}\\ \mbox{subject to}~{}\underset{h_{11},\dots,h_{lk}\in S^{\prime}}{\sup}&\sum_{l=1}^{L}g_{l}^{\prime}(x,h_{lk})-\lambda c_{k}^{\prime}(z_{lk})\leq s_{k},\quad k=1,\dots,K\end{array}$
$=\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0}{\inf}&\sum_{k=1}^{K}s_{k}\\ \mbox{subject to}&\sum_{l=1}^{L}[-g_{l}^{\prime}(x,A\cdot+b)+\chi_{S^{\prime}}(A\cdot+b)+\lambda c_{k}^{\prime}]^{\star}(0)\leq s_{k},\quad k=1,\dots,K.\end{array}$
For the new functions defined, we applied the definition of conjugate functions. Now, using the conjugate form $f^{\star}(y)=\alpha g^{\star}(y)$ of a right-scalar-multiplied function $f(x)=\alpha g(x/\alpha)$, we rewrite the above as
$=\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0}{\inf}&\sum_{k=1}^{K}s_{k}\\ \mbox{subject to}&\sum_{l=1}^{L}\alpha_{lk}[-g_{l}(x,A\cdot+b)+\chi_{S}(A\cdot+b)+\lambda c_{k}]^{\star}(0)\leq s_{k},\quad k=1,\dots,K.\end{array}$
@@
$\alpha_{lk}[(-g_{l}(x,A\cdot+b)+\chi_{S}(A\cdot+b)+\lambda c_{k})]^{\star}(0)$ $=\alpha_{lk}\underset{\omega_{lk},h_{lk},y_{lk}}{\inf}([-g_{l}]^{\star}(x,h_{lk})-b^{T}(h_{lk}+\omega_{lk})$
@@
$\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0,h_{lk},\omega_{lk},s_{k}}{\inf}&\sum_{k}^{K}s_{k}\\ \mbox{subject to}&\sum_{l=1}^{L}\alpha_{lk}([-g_{l}]^{\star}(x,h_{lk})-b^{T}(h_{lk}+\omega_{lk})\\ &+\sigma_{S}(\omega_{lk})-(A^{T}(h_{lk}+\omega_{lk}))^{T}\bar{d}_{k}+\phi(q)\lambda{\|}A^{T}(h_{lk}+\omega_{lk})/\lambda{\|}^{q}_{*}+\lambda\epsilon^{p})\leq s_{k},\\ &\hskip 56.9055pt\quad k=1,\dots,K,\quad j=1,\dots,J.\end{array}$
@@
$\begin{array}[t]{ll}\underset{\lambda\geq 0,z_{lk},\omega_{lk},s_{k}}{\inf}&\lambda\epsilon^{p}+\sum_{k}^{K}w_{k}s_{k}\\ \mbox{s.t.}&[-g_{l}]^{\star}(x,h_{lk})-b^{T}(h_{lk}+\omega_{lk})+\sigma_{S}(\omega_{lk})-(A^{T}(h_{lk}+\omega_{lk}))^{T}\bar{d}_{k}\\ &\hfill+\phi(q)\lambda\left\|A^{T}(h_{lk}+\omega_{lk})/\lambda\right\|^{q}_{*}\leq s_{k},\\ &\hfill l=1,\dots,L,\quad k=1,\dots,K,\end{array}$
@@
$\begin{array}[t]{ll}{\text{minimize}}&f(x)\\ \mbox{subject to}&\lambda\epsilon^{p}+\sum_{k}^{K}w_{k}s_{k}\leq 0\\ &[-g_{l}]^{\star}(x,h_{lk})-b^{T}(h_{lk}+\omega_{lk})+\sigma_{S}(\omega_{lk})-(A^{T}(h_{lk}+\omega_{lk}))^{T}\bar{d}_{k}\\ &\hfill+\phi(q)\lambda\left\|A^{T}(h_{lk}+\omega_{lk})/\lambda\right\|^{q}_{*}\leq s_{k},\\ &\hfill l=1,\dots,L,\quad k=1,\dots,K.\end{array}$

### v5

@@ E.4 Polyhedral reformulation
$\begin{array}[t]{ll}\underset{Av+b\in S}{\sup}&g(z,Av+b)\\ \mbox{subject to}&Dz\leq d\end{array}$
$=\begin{array}[t]{ll}\underset{Av+b\in S}{\sup}~{}\underset{h\geq 0}{\inf}~{}g(z,Av+b)-(Dv)^{T}h+d^{T}h\\ \end{array}$
@@
$=\begin{array}[t]{ll}\underset{h\geq 0,\tau}{\inf}&\tau\\ \mbox{subject to}&d^{T}h_{l}+\underset{Av+b\in S}{\sup}g_{l}(z,Av+b)-(Dv)^{T}h_{l}\leq\tau,\quad l=1,\dots,L\end{array}$
$=\begin{array}[t]{ll}\underset{h\geq 0,\omega,\tau,y,\gamma}{\inf}&\tau\\ \mbox{subject to}&d^{T}h_{l}+[-g_{l}]^{\star}(z,\omega_{l})-b^{T}(\omega_{l}+\gamma_{l})+\sigma_{S}(\gamma_{l})+\mathcal{I}_{D^{T}h_{l}=-y_{l}}\\ &\hfill+\mathcal{I}_{A^{T}(\omega_{l}+\gamma_{l})=y_{l}}\leq\tau,\quad l=1,\dots,L\end{array}$
$=\begin{array}[t]{ll}\underset{h\geq 0,\omega,\tau,\gamma}{\inf}&\tau\\ \mbox{subject to}&[-g_{l}]^{\star}(z,\omega_{l})-b^{T}(\omega_{l}+\gamma_{l})+\sigma_{S}(\gamma_{l})\leq\tau,\quad l=1,\dots,L\\ &A^{T}(\omega_{l}+\gamma_{l})=-D^{T}h_{l},\quad l=1,\dots,L,\end{array}$
@@
$\begin{array}[t]{ll}{\text{minimize}}&f(z)\\ \mbox{subject to}&[-g_{l}]^{\star}(z,\omega_{l})-b^{T}(\omega_{l}+\gamma_{l})+\sigma_{S}(\gamma_{l})\leq 0,\quad l=1,\dots,L\\ &A^{T}(\omega_{l}+\gamma_{l})=-D^{T}h_{l},\quad l=1,\dots,L.\end{array}$
@@ E.5 Mean Robust Optimization reformulation
We adapt the proof from (Wang et al.,, 2024, Appendix A.3). To simplify notation, we define $c_{k}(v_{lk})\coloneqq\|v_{lk}-\bar{d}_{k}\|^{p}-\epsilon^{p}$. We also note that $u_{lk}=Av_{lk}+b$. Then, we have
$\begin{array}[t]{ll}\underset{u_{11},\dots,u_{LK}\in S,\alpha\in\Gamma}{\sup}&\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}g_{l}(z,u_{lk})\\ \mbox{~{}~{}~{}~{}subject to}&\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}c_{k}(v_{lk})\leq 0\end{array}$
$=\begin{array}[t]{ll}\underset{u_{11},\dots,u_{LK}\in S,\alpha\in\Gamma}{\sup}&\underset{\lambda\geq 0}{\inf}~{}\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}g_{l}(z,u_{lk})-\lambda\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}c_{k}(v_{lk})\\ \end{array}$
$=\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0}{\inf}~{}\underset{u_{11},\dots,u_{LK}\in S}{\sup}&\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}g_{l}(z,u_{lk})-\lambda\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}c_{k}(v_{lk}).\\ \end{array}$
@@
$=\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0}{\inf}&\sum_{k=1}^{K}s_{k}\\ \mbox{subject to}~{}\underset{u_{11},\dots,u_{LK}\in S}{\sup}&\sum_{l=1}^{L}\alpha_{lk}(g_{l}(z,u_{lk})-\lambda c_{k}(v_{lk}))\leq s_{k},\quad k=1,\dots,K\end{array}$
$=\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0}{\inf}&\sum_{k=1}^{K}s_{k}\\ \mbox{subject to}~{}\underset{\alpha_{11}u_{11},\dots,\alpha_{LK}u_{LK}\in S^{\prime}}{\sup}&\sum_{l=1}^{L}\alpha_{lk}(g_{l}(z,(\alpha_{lk}u_{lk})/\alpha_{lk})\\ &-\lambda c_{k}((\alpha_{lk}v_{lk})/\alpha_{lk}))\leq s_{k},\quad k=1,\dots,K.\end{array}$
In the last step, we rewrote $u_{lk}=(\alpha_{lk}u_{lk})/\alpha_{lk}$, $v_{lk}=(\alpha_{lk}v_{lk})/\alpha_{lk}$, and maximized over $\alpha_{lk}u_{lk}\in S^{\prime}$, where $S^{\prime}$ is the support transformed by $\alpha_{lk}$. See Wang et al., (2024) for details. Next, we make substitutions $h_{ij}=\alpha_{lk}u_{lk}$, $y_{ij}=\alpha_{lk}v_{lk}$, and define functions $g_{l}^{\prime}(z,h_{lk})=\alpha_{lk}g_{l}(z,h_{lk}/\alpha_{lk})$, $c_{k}^{\prime}(y_{lk})=\alpha_{lk}c_{k}^{\prime}(y_{lk}/\alpha_{lk})$.
$=\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0}{\inf}&\sum_{k=1}^{K}s_{k}\\ \mbox{subject to}~{}\underset{h_{11},\dots,h_{lk}\in S^{\prime}}{\sup}&\sum_{l=1}^{L}g_{l}^{\prime}(z,h_{lk})-\lambda c_{k}^{\prime}(y_{lk})\leq s_{k},\quad k=1,\dots,K\end{array}$
$=\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0}{\inf}&\sum_{k=1}^{K}s_{k}\\ \mbox{subject to}&\sum_{l=1}^{L}[-g_{l}^{\prime}(z,A\cdot+b)+\chi_{S^{\prime}}(A\cdot+b)+\lambda c_{k}^{\prime}]^{\star}(0)\leq s_{k},\quad k=1,\dots,K.\end{array}$
For the new functions defined, we applied the definition of conjugate functions. Now, using the conjugate form $f^{\star}(y)=\alpha g^{\star}(y)$ of a right-scalar-multiplied function $f(z)=\alpha g(z/\alpha)$, we rewrite the above as
$=\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0}{\inf}&\sum_{k=1}^{K}s_{k}\\ \mbox{subject to}&\sum_{l=1}^{L}\alpha_{lk}[-g_{l}(z,A\cdot+b)+\chi_{S}(A\cdot+b)+\lambda c_{k}]^{\star}(0)\leq s_{k},\quad k=1,\dots,K.\end{array}$
@@
$\alpha_{lk}[(-g_{l}(z,A\cdot+b)+\chi_{S}(A\cdot+b)+\lambda c_{k})]^{\star}(0)$ $=\alpha_{lk}\underset{\omega_{lk},h_{lk},y_{lk}}{\inf}([-g_{l}]^{\star}(z,h_{lk})-b^{T}(h_{lk}+\omega_{lk})$
@@
$\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0,h_{lk},\omega_{lk},s_{k}}{\inf}&\sum_{k}^{K}s_{k}\\ \mbox{subject to}&\sum_{l=1}^{L}\alpha_{lk}([-g_{l}]^{\star}(z,h_{lk})-b^{T}(h_{lk}+\omega_{lk})\\ &+\sigma_{S}(\omega_{lk})-(A^{T}(h_{lk}+\omega_{lk}))^{T}\bar{d}_{k}+\phi(q)\lambda{\|}A^{T}(h_{lk}+\omega_{lk})/\lambda{\|}^{q}_{*}+\lambda\epsilon^{p})\leq s_{k},\\ &\hskip 56.9055pt\quad k=1,\dots,K,\quad j=1,\dots,J.\end{array}$
@@
$\begin{array}[t]{ll}\underset{\lambda\geq 0,h_{lk},\omega_{lk},s_{k}}{\inf}&\lambda\epsilon^{p}+\sum_{k}^{K}w_{k}s_{k}\\ \mbox{s.t.}&[-g_{l}]^{\star}(z,h_{lk})-b^{T}(h_{lk}+\omega_{lk})+\sigma_{S}(\omega_{lk})-(A^{T}(h_{lk}+\omega_{lk}))^{T}\bar{d}_{k}\\ &\hfill+\phi(q)\lambda\left\|A^{T}(h_{lk}+\omega_{lk})/\lambda\right\|^{q}_{*}\leq s_{k},\\ &\hfill l=1,\dots,L,\quad k=1,\dots,K,\end{array}$
@@
$\begin{array}[t]{ll}{\text{minimize}}&f(z)\\ \mbox{subject to}&\lambda\epsilon^{p}+\sum_{k}^{K}w_{k}s_{k}\leq 0\\ &[-g_{l}]^{\star}(z,h_{lk})-b^{T}(h_{lk}+\omega_{lk})+\sigma_{S}(\omega_{lk})-(A^{T}(h_{lk}+\omega_{lk}))^{T}\bar{d}_{k}\\ &\hfill+\phi(q)\lambda\left\|A^{T}(h_{lk}+\omega_{lk})/\lambda\right\|^{q}_{*}\leq s_{k},\\ &\hfill l=1,\dots,L,\quad k=1,\dots,K.\end{array}$